### v3

# CriSPO: Multi-Aspect Critique-Suggestion-guided Automatic Prompt Optimization for Text Generation

Abstract
Existing automatic prompt engineering methods are typically designed for discriminative tasks, where new task prompts are iteratively refined with limited feedback from a single metric reflecting a single aspect. However, these approaches are suboptimal for generative tasks, which require more nuanced guidance beyond a single numeric metric to improve the prompt and optimize multiple aspects of the generated text. To address these challenges, we propose a novel multi-aspect Critique-Suggestion-guided automatic Prompt Optimization (CriSPO) approach. CriSPO introduces a critique-suggestion module as its core component. This module spontaneously discovers aspects, and compares generated and reference texts across these aspects, providing specific suggestions for prompt modification. These clear critiques and actionable suggestions guide a receptive optimizer module to make more substantial changes, exploring a broader and more effective search space. To further improve CriSPO with multi-metric optimization, we introduce an Automatic Suffix Tuning (AST) extension to enhance the performance of task prompts across multiple metrics. We evaluate CriSPO on 4 state-of-the-art Large Language Models across 4 summarization and 5 Question Answering (QA) datasets. Extensive experiments show 3-4% ROUGE score improvement on summarization and substantial improvement of various metrics on QA.
Codehttps://github.com/amazon-science/CriSPO
1 Introduction
LLMs have emerged as powerful tools for various natural language processing tasks, including text generation (Brown et al. 2020). To fully leverage their capabilities, a critical step is to design a precise task prompt which specifies the desired behavior of the LLM to solve a task. Manual prompt engineering is often laborious, skill-intensive and sub-optimal, motivating the need for automatic prompt engineering techniques which automatically tune the task prompt.
Recent research has made notable progress in automatic prompt engineering for discriminative tasks, such as text classification (Zhou et al. 2022; Yang et al. 2023; Pryzant et al. 2023; Sordoni et al. 2024). These methods focus on optimizing task prompts for a single metric on a single aspect. The process typically involves instructing an LLM optimizer with a meta-prompt to generate new task prompts based on previously sampled task prompts and their corresponding scores. By iteratively exploring candidates and selecting the task prompt with the highest score, performance on the target metric improves over numerous iterations. However, applying these methods directly to text generation tasks, such as summarization, is sub-optimal due to challenges in obtaining effective optimization signals. Unlike classification tasks, where metrics are straightforward (eg. accuracy), automatic metrics for text generation, like ROUGE (Lin 2004), provides limited guidance for prompt refinement. For example, a lower ROUGE score may result from aspects such as mismatched length, differences in word choice due to formality, or varying writing formats, making it difficult to guide LLMs in prompt modification without fine-grained feedback targeting these individual aspects. Furthermore, evaluating text generation involves multiple metrics (Fabbri et al. 2021; Gao and Wan 2022; Elangovan et al. 2024). In addition to reference similarity, other metrics such as factual consistency, which can be assessed using metrics like AlignScore (Zha et al. 2023), are also important. Balancing or utilizing these multiple metrics is not fully addressed by existing prompt engineering methods that focus on optimizing a single metric.
To address these challenges, we introduce CriSPO, a multi-aspect Critique-Suggestion-guided automatic Prompt Optimization (CriSPO) approach. Overall, our approach employs LLMs to automatically identifies multi-aspect prompt revision suggestions, based on which prompts are automatically designed and refined (Table 8 in Appendix shows a working example of how a prompt gets revised in CriSPO). Inspired by recent self-reflection studies, where LLMs generate verbal feedback to aid in self-improvement (Gero et al. 2023; Shinn et al. 2023; Madaan et al. 2024), we designed the first key component of CriSPO: the multi-aspect critique-suggestion meta-prompt. It automatically discovers proper aspects to compare generated text with reference, write critiques of flaws (Pryzant et al. 2023) and suggestions to improve the task prompt (Figure 2 shows a word cloud of aspects identified by CriSPO, including number of words, style, and precision). Both critiques and suggestions, written in natural language, are more helpful for prompt improvement than a single ROUGE score. We then create a receptive optimizer meta-prompt that generates new prompts. In addition to conditioning on previous high-score task prompts and scores, this optimizer also reviews the past critiques and suggestions. It then generates an overall suggestion and an improved task prompt candidate in a Chain-of-Thought (CoT) (Wei et al. 2022) manner. Our approach iteratively optimizes the task prompt using LLMs similar to previous work like Optimization by PROmpting (OPRO(Yang et al. 2023), but it enriches the training signal with multi-aspect critiques and suggestions to better optimize a text generation metric. To further enhance performance by allowing the prompt to access external data, we design the task prompt template that contains placeholders for In-Context Learning (ICL) examples or retrieved contexts. The receptive optimizer meta-prompt generates these templates directly, so it can flexibly move components in task prompt for better organization.
While CriSPO offers multi-aspect guidance for optimizing text generation through critiques and suggestions, we further enhance this guidance by incorporating multiple metrics as additional teaching signals. To this end, we propose a novel Automatic Suffix Tuning (AST) extension which divides prompts into chunks conquering different metrics. Through multi-objective learning, we improve each new metric with little to no drop in existing metrics.
We test CriSPO on state-of-the-art LLMs, including Claude (Anthropic 2023, 2024), Mistral (Jiang et al. 2023) and Llama3 (MetaAI 2024), across 9 heterogeneous datasets. These include 4 summarization datasets spanning various abstractiveness, formats, and domains, as well as 5 QA datasets. Extensive experiments demonstrate that CriSPO significantly improves prompt quality and task performance over strong baselines as verified by human evaluation. We also conduct ablation study to assess the effectiveness of key ingredients.
Our contributions are summarized below:
1) We propose CriSPO, an automatic prompt engineering approach tailored for generative tasks. It discovers aspects to critique generated text and write suggestions for more effective prompt revision.
2) We conduct comprehensive experiments across multiple LLMs and datasets, demonstrating the effectiveness and robustness of our method. We show an overall 3-4% improvement on ROUGE scores with qualitative verification from human evaluation. CriSPO also obtained consistent improvements on various QA tasks.
3) We propose AST that enables prompt tuning for multiple metrics. We show that CriSPO with AST can jointly optimize AlignScore (Zha et al. 2023) for faithfulness and ROUGE for reference similarity.
2 Related Work
Figure 1: The CriSPO workflow for text generation tasks. In each iteration, a candidate task prompt $p_{t}$ is applied to $\mathcal{D}_{\texttt{trn}}$ (step 1) and evaluated using a multi-aspect critique-suggestion meta-prompt $M_{c}$ (step 2). We select top-$K$ previously sampled task prompts (step 3) and use a receptive optimizer $M_{o}$ to generate the next candidate $p_{t+1}$ (step 4). The automatic optimization loop runs multiple iterations, while the best task prompt is selected based on performance on $\mathcal{D}_{\texttt{dev}}$.
There is an increasing effort in the literature to explore gradient-free automatic prompt engineering methods with off-the-shelf LLMs. The focus of these approaches is to find a good search algorithm for better prompt candidates to solve discriminitive tasks. Earlier studies have employed conventional paraphrasing methods for prompt generation through editing phrases (Prasad et al. 2023) or back translation (Xu et al. 2022). More recently, LLMs themselves have been used to sample prompt candidates. Zhou et al. (2022) proposed Automatic Prompt Engineering (APE) which iteratively prompts an LLM to generate semantically similar variations of the locally best prompt. Pryzant et al. (2023) add verbal feedback based on error examples to propose better prompts in terms of accuracy. Concurrently, Sordoni et al. (2024) learn prompts with variational inference by considering their outputs as latent variables. Later on, Yang et al. (2023) propose OPRO to improve over them by incorporating the history of past prompts with their scores which stabilizes optimization. More structured prompts have also been explored by imposing expert-level planning (Wang et al. 2023). In a parallel thread, Fernando et al. (2023) and Guo et al. (2023) were inspired by evolutionary algorithms to perform mutation operations for prompt generation. All of the existing approaches have mostly been designed to target classification tasks using a single metric. Comparing to the existing studies, our proposed method specifically targets the unique challenges in text generation and approaches the prompt optimization problem in a multi-aspect and multi-metric fashion. For practitioners, Khattab et al. (2023) design DSPy framework to build and optimize complex LLM pipelines in a programmatic fashion. TextGrad (Yuksekgonul et al. 2024) further generalizes optimization to text beyond prompt. Our CriSPO can be used as a powerful optimizer in these frameworks.
Our approach is also inspired by recent studies on using LLMs to automatically correct its output (Pan et al. 2023; Madaan et al. 2024). Gero et al. (2023) apply multiple self-reflection steps to improve the performance of information extraction. Yan et al. (2024) use CoT to generate structured comparison and preferences for two model outputs. Shinn et al. (2023) argue the importance of the self-reflection history and propose reflexion agent to provide verbal feedback on past trials for better decision in the next trials. It is important to notice that these self-reflection studies are strictly speaking not automatic prompt engineering approaches as these studies optimize output revision rather than directly on the prompts. CriSPO, however, automatically reflects on the design of the prompt and uses these past reflections to revise the prompts.
3 Method
Problem Formulation: In a text generation task, let $\mathcal{D}_{\texttt{trn}}=\{(x_{i},y_{i})\}_{i=1\ldots n}$ be the training set, with a development set $\mathcal{D}_{\texttt{dev}}$ and a test set $\mathcal{D}_{\texttt{tst}}$. Here, $x$ represents the input data, and $y$ is the corresponding ground truth reference. A task prompt $p$ comprises instructions that, when filled with input $x$, are fed to a black-box API $\text{LLM}_{\text{task}}$ [^1] We use notations $\text{LLM}_{\text{task}},\text{LLM}_{\text{crit}},\text{LLM}_{\text{opti}}$ for clarity. Though they share the same underlying LLM unless specified otherwise. to generate a completion $\hat{y}=\text{LLM}_{\text{task}}(p,x)$. The goal is to optimize $p$ using $\mathcal{D}_{\texttt{trn}}$ and $\mathcal{D}_{\texttt{dev}}$ to identify an optimal prompt $p^{*}$ that maximizes performance on one or more evaluation metrics on $\mathcal{D}_{\texttt{tst}}$.
CriSPO Overview: CriSPO is an automatic prompt optimization algorithm designed to iteratively refine a task prompt $p$ from an initial seed prompt $p_{0}$ to the optimum: $p^{*}\leftarrow\mathcal{F}(p_{0})$. In each iteration $t$, we conduct the following steps:
  • Evaluate on $\mathcal{D}_{\texttt{trn}}$: Apply the candidate prompt $p_{t}$ on $\mathcal{D}_{\texttt{trn}}$, call $\text{LLM}_{\texttt{task}}$ to generate outputs $\{\hat{y}^{t}_{i}\}_{i=1\ldots n}$ and compute a primary metric $s_{t}$, which can be a single metric or an aggregation of multiple metrics.
  • Generate Critiques and Suggestions: Apply the multi-aspect critique-suggestion meta-prompt $M_{c}$ and call $\text{LLM}_{\texttt{crit}}$ to compare $\{\hat{y}^{t}_{i}\}_{i=1\ldots n}$ and $\{{y}^{t}_{i}\}_{i=1\ldots n}$ and generate critiques and suggestions $c_{t}$ (see Section 3.1).
  • Generate a Candidate Task Prompt: Select the top-$K$ task prompts from previous iterations based on the primary metric, and insert the corresponding $K$ triples $\{(p_{k},s_{k},c_{k})\}$ into the receptive optimizer meta-prompt $M_{o}$. Then call $\text{LLM}_{\texttt{opti}}$ to generate the next candidate prompt $p_{t+1}$ (see Section 3.2).
We evaluate the current prompt $p_{t}$ on $\mathcal{D}_{\texttt{dev}}$ and select $p^{*}$ based on the primary metric. Upon reaching the maximum number of iterations, we apply an optional AST to enhance performance on secondary metrics on $p^{*}$ (see Section 3.3). Figure 1 demonstrates the workflow of CriSPO on summarization tasks. Table 8 (in Appendix) shows a concrete working example of CriSPO.
3.1 Multi-Aspect Critiques and Suggestions
Given a prompt $p_{t}$ and its outputs $\{\hat{y}^{t}_{i}\}$ on $\mathcal{D}_{\texttt{trn}}$, we design a multi-aspect critique-suggestion meta-prompt $M_{c}$ to identify critiques – flaws of the generated outputs across multiple aspects, and suggestions – specific edits on the task prompt to rectify each flaw.
Constructive critiques with spontaneous dimension discovery:
In $M_{c}$, we first instruct $\text{LLM}_{\texttt{crit}}$ to generate several task-specific and iteration-specific aspects for a given batch of outputs from the current $p_{t}$. This approach ensures that as task prompts evolve across iterations, the focus remains on relevant aspects, addressing specific issues that arise. Figure 2 illustrates the aspects discovered during optimization. For each aspect, $M_{c}$ instructs $\text{LLM}_{\texttt{crit}}$ to generates a critique highlighting potential problems of the outputs generated with $p_{t}$ on the batch.
Figure 2: A word cloud showing the different aspects identified by CriSPO when comparing generations and references.
Multi-aspect suggestions:
In line with each critique, a corresponding suggestion is made by $\text{LLM}_{\texttt{crit}}$ to edit $p_{t}$. As opposed to Pryzant et al. (2023), we decoupled the edit suggestion module from the new prompt generation process. Rather than generating a new prompt with each suggestion, we pack a history of critiques and suggestions into the receptive optimizer for generating the next prompt, enabling more stable optimization over the infinite search space.
Our $M_{c}$ is implemented in a single CoT meta-prompt which generates, dimensions, critiques and suggestions in one single LLM call, specifically
$c_{t}=\text{LLM}_{\text{crit}}\left(M_{c},p_{t},(x_{i},y_{i},\hat{y}^{t}_{i})_{i=1\ldots n}\right).$
$M_{c}$ for different LLMs and tasks are shown in Appendix H.
Manual Automatic Prompt Engineering
0-shot 3-shot* OPRO CriSPO CriSPO 3-shot*
Dataset LLM R1 R2 RL R1 R2 RL R1 R2 RL R1 R2 RL R1 R2 RL
CNN Claude In. 37.5 12.5 22.6 40.4 14.8 24.8 39.5 14.3 24.5 40.1 15.7 26.1 42.1 17.0 27.4
Claude3 38.8 14.4 24.0 40.3 15.4 25.2 39.7 15.1 5.1 42.2 17.3 27.9 41.6 16.3 27.1
Mistral 7B 30.9 11.0 20.4 30.7 10.6 20.1 36.5 14.4 23.0 38.5 14.3 23.9 38.5 14.3 24.1
Llama3 8B 37.9 14.4 23.8 39.1 15.2 24.6# 41.5 16.3 26.5#
MeetingBank Claude In. 30.7 11.6 20.5 34.2 17.3 25.5 39.0 20.3 29.7 41.4 23.7 33.1 50.1 35.4 44.4
Claude3 31.2 14.2 22.3 37.5 22.0 29.5 41.5 21.8 32.0 47.4 32.5 40.9 58.5 46.5 54.1
Mistral 7B 26.0 11.5 18.5 31.3 14.8 22.7 33.9 15.4 24.2 39.1 19.5 29.3 35.2 16.7 26.1
Llama3 8B 31.4 14.6 22.6 40.2 22.3 31.5# 44.7 27.6 36.8#
SAMSum Claude In. 33.9 11.7 25.6 37.8 14.3 28.8 38.1 13.4 28.7 44.4 16.9 34.3 45.7 18.7 36.2
Claude3 35.8 12.7 27.0 41.1 16.6 31.3 39.0 14.7 30.1 43.4 17.1 34.3 47.2 20.8 38.2
Mistral 7B 32.0 10.2 24.1 39.5 14.1 30.3 37.9 13.6 29.0 37.6 12.4 28.4 40.0 14.2 30.8
Llama3 8B 35.7 12.3 27.1 39.3 14.7 30.0# 44.8 18.8 35.4#
ACI-Bench Claude In. 43.8 16.9 26.1 51.5 23.6 33.5 45.2 16.3 25.5 53.0 19.7 26.8 58.2 26.7 35.3
Claude3 47.3 20.3 29.3 59.1 30.1 38.6 48.8 20.1 29.5 54.0 21.4 30.3 63.1 32.5 41.0
Mistral 7B 47.8 17.7 25.4 48.4 19.2 28.1 45.1 17.0 25.2 50.2 18.2 25.6 50.3 18.7 26.2
Llama3 8B 50.5 19.8 27.7 54.2 22.0 29.3# 56.2 22.8 29.9#
Average Claude In. 36.5 13.2 23.7 41.0 17.5 28.2 40.4 16.1 27.1 44.7 19.0 30.1 49.0 24.4 35.8
Claude3 38.3 15.4 25.6 44.5 21.0 31.2 42.2 17.9 29.2 46.8 22.1 33.3 52.6 29.0 40.1
Mistral 7B 34.2 12.6 22.1 37.5 14.7 25.3 38.4 15.1 25.4 41.4 16.1 26.8 41.0 16.0 26.8
Llama3 8B 38.9 15.3 25.3 43.2 18.6 28.8 46.8 21.4 32.2
Table 1: Comparing CriSPO with manual prompts and OPRO on representative summarization benchmarks. Averaged R1/R2/RL (i.e. ROUGE-1/2/L) are reported across 3 runs. 3-shot*: 3-shot ICL with example selection. Claude3: Claude3 Sonnet. Claude In.: Claude Instant. Llama3 results marked with (#) are using Claude3 Sonnet as the optimizer. 3-shot* results are empty for Llama3 due to its limited context window. Standard deviation and SOTA results are in Appendix M. Claude3 Sonnet achieves new SOTA ROUGE-1 performance on ACI-Bench.
3.2 Receptive Prompt Optimizer
Our receptive prompt optimizer meta-prompt $M_{o}$ improves over the OPRO optimizer meta-prompt (Yang et al. 2023) by enriching its optimization trajectory $\{(p_{k},s_{k})\}$ with past critiques and suggestions $c_{k}$. Thus, ours samples candidate prompts for the next iteration conditioned on an enriched optimization trajectory:
$p_{t+1}=\text{LLM}_{\text{opti}}\left(M_{o},\{(p_{k},s_{k},c_{k})\}\right).$
Specifically, we enhance the OPRO optimizer module with the following three improvements to better utilize critiques and suggestions for achieving stronger guidance and better exploration. See Appendix I for all $M_{o}$ by LLMs and tasks.
Enriched optimization trajectory:
The critiques and suggestions generated in Section 3.1 are used in an enriched optimization trajectory to propose new prompts via an OPRO-style optimizer. Specifically, our enriched optimization trajectory includes the top-$K$ best-performing past prompts $\{p_{k}\}$, their scores $\{s_{k}\}$, critiques and suggestions $\{c_{k}\}$, sorted in the ascending order by scores. Including critiques and suggestions in the optimization trajectory allows the LLM to avoid common limitations and identify common strengths from the past prompts for stable optimization.
Chain-of-thought:
After enriching the optimization trajectory, we also apply CoT to the optimization process. Specifically, $\text{LLM}_{\texttt{opti}}$ is explicitly asked to first compare high-score prompts to low-score ones, and then elicit general ideas and learnings, and finally draft a new and better prompt. CoT further ensures the optimizer to harness collective strength from the history and identify a promising path through comparing the divergent past prompts.
Flexible task prompt template:
Instead of only tuning instruction text and fixing the input position as in existing approaches such as OPRO, CriSPO optimizes the task prompt structure using a template that can freely and naturally move around input and instruction in the prompt. It uses placeholders for the input and any external data. For example, we instruct LLM to generate example placeholder INSERT_EXAMPLES_HERE to indicate the position of ICL examples. In Retrieval-Augmented Generation (RAG) settings, we introduce a context placeholder INSERT_CONTEXT_HERE which will be replaced by the retrieved context for each question. When filled the placeholders with proper data, the task prompt clearly organized all the information to help $\text{LLM}_{\texttt{task}}$ better solve the task.
3.3 Multi-Metric Automatic Suffix Tuning
Using components in Section 3.1 and 3.2, CriSPO is ready to optimize a primary metric. To benefit from more teaching signals, e.g., completeness and faithfulness, here we extend CriSPO to multi-metric optimization by proposing a novel multi-metric learning extension named as AST.
In AST, we propose to optimize a suffix postscript $\sigma$ appended to $p^{*}$, which has already been trained on certain metrics. $p^{*}$ will remain fixed throughout the whole tuning process for a new metric to preserve most of its performance on existing metrics. $p^{*}$ is extended with an additional suffix $\sigma^{*}\leftarrow\mathcal{F}(\sigma_{0})$, which serves as a postscript to steer the LLM toward the new metric and remedy any potential regression in performance on existing metrics. Specifically, we provide both the main prompt $p^{*}$ and each suffix $\sigma_{t}$ in the meta-prompts while asking the LLM to critique or refine only the suffix. To ensure we maintain existing metrics while improving on the additional metric, we take inspirations from the balance terms of loss functions in multi-task learning (He and Choi 2021) and compute an aggregated score across the multiple metrics. Since the score of each metric is on different scales and hard to estimate before training, we propose to use the average ranking of each metric as the ultimate basis to score prompt candidates in the meta-prompt.
4 Main Experiments
4.1 Experiment Setup
Datasets
We select a diverse range of 4 summarization tasks including conventional document summarization tasks such as CNN daily mail (Hermann et al. 2015) (news headline summarization), and also conversation summarization tasks such as SAMSum (Gliwa et al. 2019), MeetingBank (Hu et al. 2023). In addition, we test on a medical-domain clinical note summarization task, ACI-Bench (Yim et al. 2023). Detailed data setup can be found in Appendix C. These tasks cover various lengths, domains and styles as summarized in Table 9. We report ROUGE-1/2/L F-measure (Lin 2004) [^2] We report additional metrics including AlignScore (Zha et al. 2023) and BertScore (Zhang et al. 2019) in Appendix M. to measure output similarity to the references.
LLMs and Baselines
We test our approach on state-of-the-art LLMs including proprietary models: Claude Instant (Anthropic 2023), Claude3 Sonnet (Anthropic 2024), and open-source LLMs: Mistral 7B (Jiang et al. 2023) and Llama3 8B (MetaAI 2024). We use the same LLM for all the 3 CriSPO modules: task inference, critique-suggestion and receptive optimization, apart from the Llama3 setup. Specific hyper-parameters with ablations are detailed in Appendix D.
Our baseline methods include manual prompts with zero/few-shot ICL. These manual prompts are carefully tuned for each task to incorporate length constraints and task guidelines, and therefore establish a high bar of performance from manual prompt engineering (Appendix J). Given there are no existing automatic prompting results for text generation, we adapted OPRO (Yang et al. 2023), a competitive established approach, on our selected tasks. We use the same hyper-parameter setup in OPRO and CriSPO for fair comparison.
4.2 Main Results
As shown in Table 1, across all the tasks and LLMs, CriSPO consistently improves over 0-shot manual prompt and OPRO baselines. Overall, there are approximately 3-4 point improvements for all LLMs. Even the strong state-of-the-art Claude3 Sonnet model can still greatly benefit from CriSPO. The consistent improvement shows CriSPO is a more effective search method than existing method (OPRO) to unlock the full potential of these LLMs, and offers an alternative solution to the more labour-intensive manual prompt engineering.
Additionally, we found examples to be helpful as adding 3-shot ICL significantly improves the performance. Owning to the versatile template in CriSPO, we can easily integrate examples and we show CriSPO 3-shot can further boost performance over CriSPO and achieves the best performance in most setups. It is also worth noticing that the vanilla CriSPO w/o ICL can match or even outperform manual prompt with 3-shot in most datasets and setups, reducing latency and cost.
4.3 Ablating Key Ingredients
Table 2 shows the ablation results of CriSPO with Claude Instant on SAMSum dataset. We observed that the three key components in our approach, including flexible template, critique and step-by-step CoT optimization, are essential for achieving optimal performance. Removing any of these components leads to a decrease in performance. Removing critique-suggestion module and CoT optimization altogether leads to a 5 point decrease, similar to OPRO performance. This indicates these two elements are essential to the success of CriSPO and flexible template is only effective when being added on top of these two elements.
Method Crit-Sugg CoT Template avg std
CriSPO 44.4 1.9
42.8 0.8
43.9 0.3
42.2 1.6
37.4 3.4
OPRO 38.1 1.3
Method Changing multi-aspect
CriSPO free multi-aspects 44.4 1.9
no multi-aspects 41.1 0.9
pre-defined multi-aspects 44.5 0.7
Table 2: Ablation studies of CriSPO on the SAMSum dataset with Claude Instant. We report average and std deviation of the ROUGE-1 F results across 3 runs.
The key novelty in our proposed novel critique-suggestion strategy in CriSPO is that it has multi-aspect: i.e. the LLM will generate multi-aspect comparison without enforcing predefined aspects. To understand the effect of the multi-aspect critique-suggestion, we provide two alternative baselines: 1. no multi-aspect: we ask LLM to compare predictions and references in general with no explicit requirement for generating critique and suggestions along multiple dimensions/aspects. This is in line with the approach adopted by Pryzant et al. (2023). 2. predefined aspects: we carefully design dimensions potentially helpful for the summarization task and include verbosity, comprehensiveness, precision and style along with their definitions (Appendix L). The no multi-aspect critique-suggestion baseline performs significantly worse, lacking critical and targeted suggestions due to its tendency to be too general. The predefined multi-aspect approach is as effective as CriSPO but we see no significant improvement from explicit definitions of dimensions. This is because the critique LLM in CriSPO is already able to identify relevant dimensions (such as completeness, verbosity etc. as in Table 8) for each iteration without explicit guidance.
4.4 Qualitative Analysis and Human Evaluation
To qualitatively compare CriSPO outputs with the baselines, we conducted human evaluation on 20 examples from the SAMSum testset. We follow the procedure from Liu et al. (2023) where the reference summaries are split into atomic content units and annotators mark them as either present or missing in the prediction summary. In total, we collected 300 annotations (100 annotations $\times$ 3 annotators). A final normalized recall score is computed with a length penalty, which indicates how similar the prediction summary is to the reference summary. In our experiment, we asked three annotators with postgraduate degrees to independently annotate the summaries with blinded setup. The inter-annotator agreement is “almost perfect” (0.8679 Fleiss kappa). We then took the majority vote, and calculated the final normalized recall score (human rating) using a de-correlated length penalty.
As shown in Table 3, CriSPO achieves the highest rating according to our human evaluation. Table 4 shows qualitative examples where prompts found by CriSPO better capture the style of the reference summaries in terms of length, what to focus on and what to skip. CriSPO outputs also look the most similar to the references, especially in terms of being as concise as the reference while covering all the key details.
Manual OPRO CriSPO
Human Rating 0.58 0.59 0.63
Table 3: Human evaluation on sampled SAMSum test.
OPRO [Best Prompt]: Generate a one to two sentence summary within the ⟨summary⟩ tags that concisely describes the key details of the conversation and any conclusions reached. INPUT_DOC
CriSPO [Best Prompt]: The text below contains a discussion expressing several key facts and events. Your concise 1-sentence summary should relate only the 2 most important pieces of information stated, without assumptions or extra context. INPUT_DOC Write the summary within ⟨summary⟩ tags.
OPRO [Example Output]: Ralph asked Andrew if he heard a Polish joke, then told a joke about sinking a Polish battleship by putting it in water. Andrew responded that the joke was terrible and so unfunny that it made his mouth dry, requiring a sip of water.
CriSPO [Example Output]: Ralph tells Andrew a Polish battleship joke that Andrew finds unfunny.
[Reference]: Ralph told Andrew a joke.
Table 4: Qualitatively comparing prompts found by OPRO and CriSPO on SAMSum. CriSPO is able to find a prompt to generate output more similar to the reference’s concise style.
4.5 Quantitative Analysis of Prompt Diversity
To verify that our design in Section 3 leads to a better exploration of the solution space, we quantitatively analyze the diversity of prompts found by CriSPO and OPRO (same hyper-parameters,  Section 4.1) on the summarization datasets. We measure 4 aggregated properties on all task prompts explored by each method during optimization: length (number of words), vocabulary size (number of unique words used), and pairwise ROUGE-L/semantic similarity. For pairwise semantic similarity, we employ Sentence Transformers (Reimers and Gurevych 2019) to obtain their embeddings and cosine distances.
As shown in Table 5, CriSPO prompts demonstrate larger variations in length and vocabulary while being less similar in lexicons and semantics, indicating its strength in exploring a larger space. We also provide a visualization of the prompts found by OPRO and CriSPO in Appendix F.
[…]
CNN
OPRO 41$\pm$6 36$\pm$5 57.5 0.93
CriSPO 149$\pm$24 96$\pm$12 50.3 0.90
MeetingBank
OPRO 31$\pm$5 28$\pm$4 44.9 0.84
CriSPO 216$\pm$41 135$\pm$19 39.7 0.80
SAMSum
OPRO 34$\pm$6 30$\pm$5 57.0 0.94
CriSPO 172$\pm$22 112$\pm$12 46.0 0.88
ACI-Bench
OPRO 58$\pm$11 46$\pm$8 62.7 0.95
CriSPO 247$\pm$40 117$\pm$13 54.3 0.93
Table 5: Prompt diversity on 4 summarization datasets.
5 Extension with Multi-Metric Optimization
AST Setup
In this experiment, we extend CriSPO with our proposed AST to optimize multiple metrics simultaneously. Specifically, we take the best prompts optimized for ROUGE-1 F-measure from CriSPO with Claude Instant as the seed main prompt $p^{*}$. We employ AST to optimize AlignScore (Zha et al. 2023) starting from a simple seed suffix $\sigma_{0}$: “Every word of your summary must be faithful to the input/conversation” across all datasets. The AlignScore between the input text and the output summary is used as a signal reflecting the faithfulness. With regard to baselines, we report the initial performance in ROUGE-1 F-measure and AlignScore of the seed main prompt w/ and w/o the seed suffix. We also provide a strong baseline to tune both the main prompt and its suffix together (full tuning) rather than only the suffix in AST.
Results
The results for multi-metric optimization are presented in Table 6. On all datasets, our AST is able to optimize the new metric AlignScore with a negligible or zero regression on the existing metric ROUGE, meaning that AST can reduce LLM hallucination while maintaining relevancy in the output. In particular, AST dramatically improves AlignScore by $11.7$ points on CNN. Across tasks, AST is the most effective approach to improve AlignScore while maintaining ROUGE. Among all methods, AST is the only one that brings consistent improvement on AlignScore for every task, and achieves the best average overall improvement (by $4.3$). The main prompt w/ suffix seed prompt slightly improves AlignScore (by $1.2$) and the full-tuning baseline only meaningfully improves AlignScore on CNN and the overall improvement is marginal (by $0.7$). The superiority of AST shows that it can robustly optimize multiple metrics across various domains.
Seed CriSPO: $\mathcal{F}(\cdot)$
main w/ suffix full w/ AST
Dataset $p^{*}$ $p^{*}+\sigma_{0}$ $\mathcal{F}(p^{*}+\sigma_{0})$ $p^{*}+\mathcal{F}(\sigma_{0})$
CNN
ROUGE-1 40.7 40.6 40.6 40.4
AlignScore 66.5 69.5($\uparrow$3.0) 69.6($\uparrow$3.1) 78.1($\uparrow$11.7)
MeetingBank
ROUGE-1 39.6 39.9 39.4 39.7
AlignScore 43.6 43.7 43.8 44.4($\uparrow$0.9)
SAMSum
ROUGE-1 45.5 45.9 45.8 45.1
AlignScore 87.2 86.6($\downarrow$0.6) 86.6($\downarrow$0.6) 88.6($\uparrow$1.4)
ACI-Bench
ROUGE-1 54.4 55.2($\uparrow$0.8) 54.5 54.3
AlignScore 66.7 69.0($\uparrow$2.3) 66.5 70.0($\uparrow$3.4)
Average
ROUGE-1 45.1 45.4 45.1 44.9
AlignScore 66.0 67.2($\uparrow$1.2) 66.7($\uparrow$0.7) 70.3($\uparrow$4.3)
Table 6: Claude Instant multi-metric results on summarization tasks. In seed block, main and w/ suffix refers to the ROUGE-1-optimized prompt $p^{*}$ and its concatenation with the manual suffix $\sigma_{0}$ respectively. In CriSPO block, full and w/ AST refers to the full-prompt tuning baseline and AST in Section 3.3 respectively. We show in the parentheses the absolute differences with main $p^{*}$ only when $>0.5$. Here we use CriSPO $\mathcal{F}(\cdot)$ to optimize an aggregation of two metrics.
6 Generalization to Other Tasks
To confirm its generalizability, in this section, we apply CriSPO to extractive, abstractive and multi-choice QA tasks.
Datasets
We benchmark CriSPO on 5 commonly used QA datasets, including 1) Wikipedia-based QA: Natural Questions (Kwiatkowski et al. 2019), TriviaQA (Joshi et al. 2017), Squad (Rajpurkar et al. 2016) 2) story-based abstractive reading comprehension: NarrativeQA (Kočiskỳ et al. 2018) and 3) medical domain multiple-choice QA: MedMCQA (Pal, Umapathi, and Sankarasubbu 2022) . For Natural Questions and TrivialQA, we also incorporate the RAG setup to optimize the prompt template with inserted pre-retrieved contexts from each dataset. We retrieved the Wikipedia pages following Izacard and Grave (2021). For NarrativeQA, we use summaries as contexts. For MedMCQA, we cast it to text generation by eliciting reasoning before the final answer. Following the conventions, we report Exact Match for Natural Questions and TriviaQA, F1 for Squad, ROUGE-L for NarrativeQA, accuracy for MedMCQA. For efficiency, we only used a small fraction of the train and dev set for the experiments. The specific data settings are listed in Appendix C.
Results
Similar to summarization tasks, we observe CriSPO significantly outperforms the manual prompt and OPRO baseline in various QA datasets as shown in Table 7. For NarrativeQA, CriSPO brings massive improvement ($+10$ ROUGE-L) compared with baselines, achieving the new SOTA performance. For Natural Questions and TrivialQA, CriSPO has no issue incorporating the RAG setup and achieving consistent improvement over the manual prompt and OPRO. Surprisingly, CriSPO even outperforms OPRO on MedMCQA despite it is not designed for classification tasks.
Manual Automatic Prompt Engineering
Task Claude 0-shot 64* OPRO CriSPO CriSPO 64*
NQ Instant 34.0 33.4 8.0 36.5 37.8
Sonnet 26.6 32.0 6.7 38.3 38.7
T-QA Instant 58.6 59.2 53.7 66.3 67.5
Sonnet 58.4 65.0 41.8 70.6 72.1
0-shot 5* OPRO CriSPO CriSPO 5*
Squad Instant 79.5 82.5 78.5 87.8 89.4
Sonnet 76.1 83.2 76.4 85.3 87.9
NarQA Instant 64.2 67.0 59.4 75.1 76.1
Sonnet 64.0 66.7 58.6 76.2 75.2
Med- Instant 49.2 53.8 50.5 52.3 54.4
MCQA Sonnet 49.8 54.4 57.7 57.9 57.4
Table 7: Comparing CriSPO with manual prompts and competitive automatic prompt engineering baseline OPRO on representative QA benchmarks. We report Exact matching for NQ (Natural Questions) and TQA (TrivialQA). We report F1 for Squad, Rouge-L for NarQA (NarrativeQA), and accuracy for MedMCQA. $k$*: $k$-shot ICL with example selection. Standard deviations can be found in Table 17.
7 Conclusion
In this paper, we tackle the challenging problem of automatic prompt engineering for text generation. We propose CriSPO, a multi-aspect critique-suggestion guided optimizer augmented with enriched trajectory, CoT and flexible template. Our experiments show multi-aspect critique-suggestion is critical for finding good task prompts. Overall, CriSPO achieves 3-4% ROUGE score improvement and 4-5% human rating increase compared to baseline methods for summarization, and significant improvement for QA. We also show that CriSPO can effectively optimize multiple metrics through a novel suffix tuning extension AST, and incorporate ICL and RAG with flexible prompt templates. Ablation studies confirm the effectiveness of all CriSPO components. Human evaluation and quantitative analysis show CriSPO encourages more effective prompt exploration and the optimized prompts can better capture task requirements.
Limitations
The list of LLMs in our experiments is meant to be representative rather than exhaustive. We recognize that supervised fine-tuning can outperform prompt engineering on certain metrics. We also acknowledge the ongoing research on the limitations of automatic evaluation metrics for text generation. In addition, CriSPO could be costly in LLM API tokens especially with long input. Finally, while our experiments focus on summarization and QA, CriSPO can be adaptable to other text generation tasks, which we leave for future research. See Appendix B for a detailed limitations discussion.
Appendix
Flow Component Text
Prompt v1
Here is an input: INSERT_INPUT_HERE Write a summary within ⟨summary⟩ tags for it.
Output
score: 33.3
Tegan tells Valentia that Paul’s brother sent her a friend request on social media. She is unsure how to respond to the request. Valentia advises Tegan to ignore the friend request, as she does not want Tegan to go down the same difficult road she had with Paul. Tegan agrees with Valentia’s perspective.
Critique - Number of words: The predicted summaries tended to be longer with more details while the reference summaries were shorter and more concise. - Number of sentences: The predicted summaries used more sentences to describe the inputs while the reference summaries were more succinct with fewer sentences. - Precision: Some details in the predicted summaries were not important and not mentioned in the reference summaries. - Recall: Some key details highlighted in the reference summaries were missing from the predicted summaries.
Suggestion - Specifying the expected length of the summary (e.g. 1-2 sentences) - Emphasizing to only include the most important/key details - Indicating which details should be prioritized or omitted
Prompt v2
For the given text, write a 1-2 sentence summary within ⟨summary⟩ tags that highlights the most important details. Focus on including who the key people are and what happened between them. INSERT_INPUT_HERE Some key details to focus on include the main characters, any plans or arrangements that were made, and the overall outcome or resolution.
Output
score: 42.1
Tegan tells Valentia that Paul’s brother sent her a friend request on social media, though she and Paul had previously broken up. Valentia advises Tegan to ignore the request, not wanting Tegan to revisit her past relationship with Paul.
Critique
- Number of words: Predicted summaries tend to be longer than reference summaries. - Number of sentences: Predicted summaries use multiple sentences while references use 1-2 sentences. - Precision: Predicted summaries include more contextual details while references focus only on key people and events. - Recall: Predicted summaries cover more information from the input while references only highlight the most important details.
Suggestion
- Suggest using less than 10 words where possible - Emphasize concise 1-2 sentence format over detail - Specify to focus only on key events/outcomes, not context
CoT
Suggestion
Comparing the high-score and low-score instructions, here are some suggestions that could improve them: ⟨suggestion⟩ Specify the desired length or range for the summaries (e.g., 10 words and 1-2 sentences).⟨/$s$uggestion⟩ Specify to focus on key events and specify which details ⟨/suggestion⟩ ⟨suggestion⟩ Specify the output should not contain unnessary context ⟨/suggestion⟩
Improved
Prompt
Read the dialogue provided in INSERT_INPUT_HERE and identify the key events between characters and outcomes. Then write a 1-2 sentence summary within ⟨summary⟩ tags that concisely captures these important plot points, such as who will borrow a dress or who has an interview, while keeping within 10 words where possible. Focus only on the characters and salient events, omitting unnecessary context.
Improved
Output
score: 75.6
Tegan receives a friend request from Paul’s brother and Valentia advises her to ignore it due to past issues.
Reference
Tegan has received a friend request from Paul’s brother. Valentia advised her not to accept it.
Table 8: A working example for the CriSPO framework for SAMSum summarization with Claude3 Sonnet on one train data point. Yellow highlights the generation from the multi-aspect critique-suggestion module. Green highlights the generation from the receptive optimizer module. The score is ROUGE-1 F for the data point.
Appendix A A Complete Working Example
Table 8 shows a full working example of CriSPO.
Appendix B Limitations
Minor prompt adaptation for different LLMs.
Different LLMs have varying context length limits, preferred input/output formats, etc. Therefore, our approach still requires some manual adaptation to different LLMs. However, the manual effort is significantly less compared to manually tuning task-specific prompts, because: 1) once tuned, the crit and opti prompts can be reused for different tasks, and 2) the tuning should mainly focus on formatting input/output, and adjusting the number of examples to fit the context length, which are straightforward following the LLM documentation.
Evaluation metrics.
Evaluating text generation is a challenging problem in itself. For summarization, our work focuses on ROUGE scores to quantify the similarity between generated and reference texts, and AlignScore to evaluate the factuality of the generated text. We also conducted human evaluation to verify our findings. (Augenstein et al. 2023) calls out that current factuality evaluations are not reliable. (Elangovan et al. 2024) highlights the challenges of conducting human evaluation in LLM era. However, we acknowledge that these evaluations are still limited, while designing better evaluation metrics is beyond the scope of this paper.
Comparing to SOTA SFT models.
We would like to emphasize that CriSPO is not designed to outperform the state-of-the-art gradient-based supervised fine-tuning (SFT) models. For some datasets, our approach still falls short compared to SOTA SFT models. Prompt tuning is a discrete optimization process with noisy directional signals on top of a limited number of prompt tokens, compared to supervised fine-tuning, which uses continuous gradient descent on much larger datasets to optimize much more parameters. Therefore, it is usually harder to match the performance of SFT.
Comparison between LLMs.
Our benchmark on various LLMs is designed to demonstrate that CriSPO is compatible with a wide range of both proprietary and open-weight (lightweight) LLMs. The list of LLMs is meant to be representative instead of exhaustive. We acknowledge the existence of more powerful LLMs from each family that may push the performance even higher, which we leave for future work.
Generalization beyond summarization and QA.
In our experiments, we mainly focused on summarization and question answering tasks. However, our proposed approach is general and can adapt to various text generation tasks, since the LLM-based critique-suggestion model only takes generated text and reference text as input and can spontaneously compare them along relevant dimensions. Our framework can potentially benefit classification tasks other than MedMCQA if they provide “explanation” or “reasoning” to each label.
Cost
Despite CriSPO has been optimized to use a relatively smaller number of candidates than existing methods per each step, it still requires a full evaluation of the candidates on the sampled training set of 50-200 examples, which costs significant amounts of LLM API tokens and time considering the optimization runs for 100 iterations. Especially when the inputs involve long contexts (e.g., RAG) and/or the training set is large. In our RAG settings, the optimization takes up to 2 days to finish.
Dataset Description Input Output
CNN/DailyMail News article headline generation. 773 58
MeetingBank City council meeting (long conversation) summarization 3095 66
SAMSum Messenger-like (short) conversation summarization 127 23
ACI-Bench Docter-patient (long) conversation medical note generation 1372 476
Natural Questions Open-domain QA using RAG on Wikipedia 20009.2 2.2
TriviaQA Open-domain QA using RAG on Wikipedia 20016.4 2.8
SQuAD Reading comprehension on Wikipedia 149.7 3.4
NarrativeQA Story reading comprehension 653.6 5.0
MedMCQA Multiple-choice QA in medical domain 38.0 100.6
Table 9: Dataset description and the number of words in input and output. For QA datasets, the input includes both contexts and the question.
(a) The effect of number of examples in the Critique-Suggestion meta-prompt
(b) The effect of number of the history prompts in the receptive optimizer meta-prompt
(c) The effect of the train/dev sample size
(d) The effect of the number of iterations
Figure 3: Ablation studies with Claude Instant on SAMSum. For each setup, we report the mean ROUGE1 F1 and standard deviation across three runs. For (c), we change the random seed to select different samples across the three runs.
(a) CNN
(b) MeetingBank
(c) SAMSum
(d) ACI-Bench
Figure 4: Visualization of prompt diversity on 4 summarization datasets for OPRO in red $\bullet$ and CriSPO in blue $\times$.
Appendix C Dataset Setting
For CNN, SAMSum, MeetingBank, MedMCQA, Narrative QA and SQUAD, we use the HuggingFace datasets repository. For ACI-Bench, we use the data from Task B at ACL ClinicalNLP MEDIQAChat shared task 2023 in the Acibench dataset [^3] https://github.com/abachaa/MEDIQA-Chat-2023 (Yim et al. 2023). For Natural Questions, we follow the data preparation in FiD [^4] https://github.com/facebookresearch/FiD (Izacard and Grave 2021).
Our experiments are conducted with sampled train and dev set. For ACI-Bench, we used the full training (67), development (20) and test set (40). For other summarization tasks, we randomly selected 500 samples from the full test set as our test set. To show the efficiency of our approach, we used a small fraction of the train and development set. For CNN, we sampled 100 training samples as our training set, and 100 development samples as our development set. For other tasks, we randomly sampled 50 training samples as our training set, and 50 development samples as our development set.
For NQ and TQA, we randomly sample 200/200/500 examples for training/development/test set. Each example has 100 context paragraphs from Wikipedia and each paragraph has 100 words following Izacard and Grave (2021). We use only the top 20 context paragraphs in our experiments because of the high inference cost for long text.
For NarrativeQA and MedMCQA, we randomly sample 100/100/500 for training/development/test set respectively. For Squad, we sample 50/50/500 for training/development/test set respectively.
Appendix D CriSPO Settings for Different LLMs
Claude Settings: In suggestion-critique meta-prompt, we pass 10 randomly selected examples for the LLM to provide critique. In optimizer meta-prompt, we use 10 history task prompts with their critiques, suggestions and scores. We add 2 input/output examples in the optimizer prompt.
Mistral Settings: Mistral has a shorter context window. Therefore, we adjust the settings. We reduce the task prompt history to 1 in optimizer meta-prompt. On MeetingBank dataset, we truncate the input document to 3500 words, and do not provide the input (only use generated text and reference) in the critique-suggestion meta-prompt.
Llama3 Settings: The context window of Llama3 is insufficient to fit a few examples and generate meaningful critique-suggestions. Therefore, we use Claude3 Sonnet as the critique-suggestion LLM and the receptive optimizer LLM. Llama3 is used only as the task LLM.
For all experiments, we set the temperature of the meta-prompt LLMs used for the optimization to be 1.0 to encourage diversity, and we set the scorer LLM’s temperature to be 0 which gives more stable results when the LLM is performing inference with the task prompt. We use the same LLM as meta-prompt and task prompt except for the Llama 3.
The initial prompts are generic and naive prompt. For example, for summarization, the starting prompt is “Generate a summary for the input text”. For QA, the starting point is “Answer the question using the context provided”. We also follow the original OPRO paper (Yang et al. 2023) to sample k prompt candidates at each step for CriSPO. We set k = 3 in our experiments. For efficiency, we also run dev set evaluation every 5 steps rather than on every step. We end the optimization process when we reach 100 steps.
Appendix E Hyper-parameter Search
We conducted experiments as shown in Figure 3 to assess the effect of different hyper-parameters. We show the performance increases as we increase the number of examples in the critique-suggestion meta-prompt and 10 examples have a significant jump of performance compared with 1 or 5 examples in the prompt. In line with Yang et al. (2023), we also found the history of prompts is helpful, where we see significant improvement when we increase from 1 prompt history to 10 prompt history, but no significant difference as we further increase the history to 20. As to the sample size, the performance grows at the beginning when we increase the size from 10 to 50, and plateaus moving from 50 to 100. We choose 50 sample for most of our experiments as it is a sufficiently representative sample to achieve good performance with relatively lower latency. We observed larger variations and generally lower performance when the number of iterations are below 100, but further iterations above 100 also does not further improve results as it is most likely that we already found the best prompt within 100 iterations. In conclusion, the most optimal combination of these hyper-parameters are: 10 examples in the critique prompt, 10 or 20 history prompts, with 50 train/dev set, with 100 iterations.
Appendix F Visualization of Prompt Diversity
To verify that our design in Sec. 3 leads to a better exploration of the solution space, we examine the distributions of prompts found by CriSPO and OPRO in an embedding space. We first encode their Claude Instant prompts on the 4 summarization datasets using the all-MiniLM-L6-v2 model from Sentence Transformers (Reimers and Gurevych 2019). Then, we perform t-SNE visualization (Van der Maaten and Hinton 2008) to their embeddings in a two-dimensional map.
As illustrated in Figure 4, CriSPO produces more diverse prompts than OPRO on all of the 4 datasets. The distribution of OPRO prompts is more centralized, indicating that OPRO prompts are homogeneous in semantics, possibly the “semantically similar paraphrases” (Yang et al. 2023). However, CriSPO distribution is more divergent, with prompts spread out over a wider range. This visualization suggests that prompts tuned by CriSPO are semantically more dispersed and versatile, which expand the exploration beyond paraphrasing and directionless Monte-Carlo search.
Appendix G Experiments on Other Tasks
We conduct experiments on other natural language generation (NLG) tasks – DailyDialog (Li et al. 2017) and WebNLG (Gardent et al. 2017). The results are shown on table 10.
Method DailyDialog (R-L) WebNLG (BLEU)
Manual 0-shot 12.6 31.3
Manual 3-shot 17.1 34.8
OPRO 13.3 33.3
CriSPO 17.4 44.3
Table 10: Results on DailyDialog and WebNLG.
Appendix H Multi-Aspect Critique-Suggestion Meta-Prompt
H.1 Claude for Summarization
In a summarization task, a writer is given an input text to write a summary following an instruction.
<instruction>{instruction}</instruction>
<examples>
<example>
<input>
{document}
</input>
<predicted_summary>
{predicted_summary}
</predicted_summary>
<reference_summary>
{reference_summary}
</reference_summary>
</example>
</examples>
Write a general and helpful critique in <critique> XML tags to improve the instruction such that the predicted summaries are as close to references as possible.
1. Come up with several dimensions to compare its predicted summaries and reference summaries, e.g., number of words, number of sentences, style, precision, recall, etc.
2. List the difference predicted summaries and references on each dimension.
3. Identify specific phrases in the instruction that could have gotten these predicted summaries different with references on each dimension.
4. Suggest specific action items that are general to all examples and helpful to improve the instruction.
H.2 Mistral for Summarization
In a summarization task, a writer is given an input text to write a summary following an instruction.
INSTRUCTION:
{instruction}
Here are a few examples using the instruction.
EXAMPLE {id}
INPUT:
{document}
PREDICTED_SUMMARY:
{predicted_summary}
REFERENCE_SUMMARY:
{reference_summary}
Write a general and helpful critique to improve the instruction such that the predicted summaries are as close to references as possible.
1. Come up with several dimensions to compare its predicted summaries and reference summaries, e.g., number of words, number of sentences, style, precision, recall, etc.
2. List the difference predicted summaries and references on each dimension.
3. Identify specific phrases in the instruction that could have gotten these predicted summaries different with references on each dimension.
4. Suggest specific action items that are general to all examples and helpful to improve the instruction.
H.3 Claude for RAG
In a question-answering task, question and context are provided and the answer needs to be generated.
<instruction>{instruction}</instruction>
<examples>
<example>
<question>
{question}
</question>
{context}
<generated_answer>
{generated_answer}
</generated_answer>
<gold_answer>
{gold_answer}
</gold_answer>
</example>
</examples>
Write a general and helpful critique in <critique> XML tags to improve the instruction such that the generated answer are the same as gold answer.
1. Come up with several dimensions to compare its generated and gold answer, e.g., number of words, style, precision, recall, etc.
2. List the difference between generated and gold answer on each dimension.
3. Identify specific phrases in the instruction that could have gotten these generated answer different with gold one on each dimension.
4. Suggest specific action items that are general to all examples and helpful to improve the instruction.
Appendix I Receptive Optimizer Meta-Prompt
I.1 Claude for Summarization
Your task is to optimize the instruction for a summarization task, where a writer is given an input text to write its summary following your instruction.
Below are some examples:
<example>
<instruction>?</instruction>
<input>
{article}
</input>
<summary>
{summary}
</summary>
</example>
Below are some previous instructions with their scores and critiques.
<rated_instruction>
<instruction>{instruction}</instruction>
<score>{score}</score>
<critique>
{critique}
</critique>
</rated_instruction>
Generate an instruction that is different from all the instructions above, and has a higher score than all the instructions above.
It should be concise, effective, and generally applicable to all examples above.
Draft your new instruction step by step:
1. Compare high-score instructions to low-score ones, identify what suggestions could have improved them. List them in <suggestion> tags.
2. Apply the suggestions and draft a new instruction aiming for a higher score.
3. Be creative and vary the wording, paraphrase, position of INSERT_INPUT_HERE and INSERT_EXAMPLES_HERE, phrase order, grammar, sentence order and etc.
4. Write your final new instruction in <instruction> tags.
I.2 Mistral for Summarization
Your task is to optimize the instruction for a summarization task, where a writer is given an input text to write its summary following your instruction.
Below are some examples:
EXAMPLE {id}
INPUT:
{article}
TARGET_SUMMARY:
{summary}
Below are some previous instructions with their scores and critiques.
INSTRUCTION:
{instruction}
SCORE:
{score}
CRITIQUE:
{critique}
Generate an instruction that is different from all the instructions above, and has a higher score than all the instructions above.
It should be concise, effective, and generally applicable to all examples above.
Draft your new instruction step by step:
1. Compare high-score instructions to low-score ones, identify what suggestions could have improved them. Write down your suggestions first.
2. Apply the suggestions and draft a new instruction aiming for a higher score.
3. Be creative and vary the wording, paraphrase, position of <INSERT_INPUT_HERE> and <INSERT_EXAMPLES_HERE>, phrase order, grammar, sentence order and etc.
4. Write your final new instruction in <instruction></instruction> tags.
5. In your final prompt, you must use <INSERT_INPUT_HERE> only once and use it in a separate line.
6. In your final prompt, you must use <INSERT_EXAMPLES_HERE> only once and use it in a separate line.
I.3 Claude for RAG
Your task is to optimize the instruction for a question-answering task, where the question and context are provided.
Below are some examples:
<example>
<instruction>?</instruction>
<question>
{question}
</question>
{context}
<answer>
{answer}
</answer>
</example>
Below are some previous instructions with their scores and critiques.
<rated_instruction>
<instruction>{instruction}</instruction>
<score>{score}</score>
<critique>
{critique}
</critique>
</rated_instruction>
Generate an instruction that is different from all the instructions above, and has a higher score than all the instructions above.
It should be concise, effective, and generally applicable to all examples above.
Draft your new instruction step by step:
1. Compare high-score instructions to low-score ones, identify what suggestions could have improved them. List them in <suggestion> tags.
2. Apply the suggestions and draft a new instruction aiming for a higher score.
3. Be creative and vary the wording, paraphrase, position of "{question_placeholder}", "{context_placeholder}", phrase order, grammar, sentence order, which specific examples to give, etc.
4. Write your final new instruction in <instruction> tags.
Appendix J Manual Prompts
We present the manual prompts for the summarization experiments with the Claude instant model. INSERT_INPUT_HERE in each prompt indicates the position where we will insert the input text. INSERT_EXAMPLES_HERE indicates the position where we will insert few-shot examples. Each example is in the format of
<examples>
<input> ...<input>
<summary> ... <summary>
</examples>
For the few-shot setup, we first encode inputs with BERT embeddings (Devlin et al. 2019), then retrieve their most similar examples from the train set according to the cosine similarity (Liu et al. 2022).
J.1 Zero-shot CNN
Here is an input CNN news document:
INSERT_INPUT_HERE
Please write a headline summary between around 50 to 100 words within <summary> tags.
J.2 Few-shot CNN
Write a headline summary between around 50 to 100 words for the CNN news document. Here are example input documents and example output summaries
INSERT_EXAMPLES_HERE
Here is an input CNN news document:
INSERT_INPUT_HERE
Please write a headline summary between around 50 to 100 words within <summary> tags.
J.3 Zero-shot SAMSum
Here is an input conversation:
INSERT_INPUT_HERE
Please write a summary for the input conversation within <summary> tags. The summary should (1) be rather short with 20 to 50 words, (2) extract important pieces of information, (3) include names of interlocutors, (4) be written in the third person.
J.4 Few-shot SAMSum
Write a summary within <summary> tags for the input conversation. Here are example input conversations and example output summaries
INSERT_EXAMPLES_HERE
Here is the input conversation:
INSERT_INPUT_HERE
Following the examples, please write a summary for the input conversation within <summary> tags. The summary should (1) be rather short with 20 to 50 words, (2) extract important pieces of information, (3) include names of interlocutors, (4) be written in the third person.
J.5 Zero-shot MeetingBank
Here is an input conversation from city council meeting:
INSERT_INPUT_HERE
Please write a summary of the discussion with around 60 to 150 words within <summary> tags.
J.6 Few-shot MeetingBank
Write a summary for the input city council meeting. Here are example input meeting conversations and example output summaries
INSERT_EXAMPLES_HERE
Here is an input conversation from a city council meeting:
INSERT_INPUT_HERE
Following the examples, please write a summary of the discussion from the input conversation with around 60 to 150 words within <summary> tags.
J.7 Zero-shot ACI-Bench
Here is an input conversation of a clinical visit:
INSERT_INPUT_HERE
Please write a detailed clinical note summary for the input conversation within <summary> tags.
J.8 Few-shot ACI-Bench
Write a clinical note summary within <summary> tags for the input conversation of a clinical visit. Here are example input conversations and example output summaries
INSERT_EXAMPLES_HERE
Here is the input conversation:
INSERT_INPUT_HERE
Following the examples, please write a clinical note summary for the input conversation within <summary> tags.
J.9 Manual Prompt Tuning
While it is not possible to exhaust all prompt variations with manual prompt engineering, we experimented with several iterations of manual prompts and presented the best prompt results. Below, we show that our tuned zero-shot manual prompts (ours) significantly outperform zero-shot naive prompts (”Write a summary for the input text”), and the results from our manual prompts can be regarded as a reasonable baseline from human prompt engineering.
CNN MBank SAMSum ACI-Bench
Naive 34.8 29.7 29.9 34.3
Ours 37.5 30.7 33.9 43.8
Table 11: Comparing our zero-shot manually tuned prompts (ours) with naive prompts on summarization tasks with Claude Instant. MBank= MeetingBank.
Appendix K Best QA Prompts Found using CriSPO (Claude Instant)
K.1 Natural Questions
Consider INSERT_QUESTION_HERE and all provided INSERT_CONTEXT_HERE. Write a concise answer in <answer> tags focusing only on the single most important attribute implied across contexts. Then compare your answer to the gold below through reasoning: cite how your intended meaning matches theirs on attributes like level of precision/detail implied jointly by contexts. It is acceptable for your answer to have less context than the gold if the meaning remains clear, like using a single word versus a phrase. Explain any differences using specific examples from contexts. Answers should be as concise as possible while still encompassing implications as fully as contexts allow.
K.2 TriviaQA
Read the question and contexts carefully. Extract the key detail(s) directly answering the question from the most relevant context(s). Write your response in <answer> tags matching the style and level of detail of the example gold answers. Consider using a single word, number, or short phrase if that fully answers the question precisely. Compare your answer to the examples, considering alternatives suggested in the contexts and relationships between entities. Aim for consistency with the gold answers in terms of words used, precision, and completeness of specification.
INSERT_CONTEXT_HERE
INSERT_QUESTION_HERE
K.3 MedMCQA
QUESTION_PLACEHOLDER Provide your answer, and comprehensively reason through it by: referencing authoritative medical sources, accounting for all relevant context in the question, logically laying out your reasoning steps, and addressing any applicable exceptions or nuances. Your response should demonstrate a rigorous application of established medical knowledge.
Chose an option and write it in <answer> XML tags
K.4 NarrativeQA
Provide a focused, concise answer in the form of a 1-3 word phrase or brief quote, enclosed in <answer> tags. Capture all key details directly relevant to fully addressing the question, while excluding extraneous background information or repetition of context details. If a short quote from the context directly and precisely answers the question in a maximally concise manner, use the quote verbatim. Otherwise, paraphrase the essential information as succinctly as possible. The goal is a clear, to-the-point response that comprehensively answers the core of the question without omitting crucial details or including unnecessary information.
CONTEXT_PLACEHOLDER
QUESTION_PLACEHOLDER
K.5 Squad
INSERT_CONTEXT_HERE
INSERT_QUESTION_HERE
Your task is to answer the question as concisely as possible using only the minimum information explicitly asked for. Carefully examine the question to understand exactly what specific detail is being requested, then scan the context to extract only that precise piece of information to satisfy the question - no more and no less. Avoid including any additional context, descriptors or embellishments beyond the single term or brief phrase strictly necessary to directly answer what is asked. Refer to the examples, where "pub landlord" and "French alone is the official language" are the minimum possible responses. Do not exceed these examples in length or level of detail. Write only the clearest, most succinct answer in <answer> tags.
Appendix L Ablation Study Prompts
Pre-defined multi-aspect critique-suggestion meta-prompt:
- Verbosity and length: compare the level of details and the length between prediction and reference summaries
- Comprehensiveness: compare whether the prediction covers all the information from the reference summaries
- Precision: compare whether the information from the prediction summaries are present in the reference summaries.
- Style: compare the formatting, formality, word choices, sentence structures etc.
Appendix M Full Metrics for Summarization
We report the average and standard deviation from 3 runs for Rouge1 (Table 12), Rouge2 (Table 13), RougeL (Table 14), BertScore (Table 15) and AlignScore (Table 16).
Manual Automatic Prompt Engineering
Dataset LLM 0-shot 3-shot* OPRO CriSPO CriSPO 3-shot*
CNN Claude Instant 37.5 40.4 39.5 (±0.4) 40.1 (±0.5) 42.1 (±0.6)
SOTA: 48.2 Claude3 Sonnet 38.8 40.3 39.7 (±0.6) 42.2 (±0.9) 41.6 (±1.0)
 (Mu and Lim 2022) Mistral 7B 30.9 30.7 36.5 (±1.8) 38.5 (±1.7) 38.5 (±1.0)
Llama3 8B 37.9 39.1 (±0.3)# 41.5 (±0.7)#
MeetingBank Claude Instant 30.7 34.2 39.0 (±6.1) 41.4 (±2.4) 50.1 (±0.6)
SOTA: 70.3 Claude3 Sonnet 31.2 37.5 41.5 (±2.2) 47.4 (±1.7) 58.5 (±1.3)
 (Hu et al. 2023) Mistral 7B 26.0 31.3 33.9 (±3.7) 39.1 (±4.8) 35.2 (±0.7)
Llama3 8B 31.4 40.2 (±3.0)# 44.7 (±0.8)#
SAMSum Claude Instant 33.9 37.8 38.1 (±1.3) 44.4 (±1.9) 45.8 (±0.4)
SOTA: 55.3 Claude3 Sonnet 35.8 41.1 39.0 (±1.4) 43.4 (±2.1) 47.2 (±0.3)
 (Wang, Liu, and Chen 2023) Mistral 7B 32.0 39.5 37.9 (±0.8) 37.6 (±3.4) 40.0 (±1.0)
Llama3 8B 35.7 39.3 (±0.6)# 44.8 (±3.4)#
ACI-Bench Claude Instant 43.9 51.5 45.2 (±0.2) 53.0 (±0.4) 58.2 (±1.8)
SOTA: 53.5 Claude3 Sonnet 47.3 59.1 48.8 (±1.9) 54.0 (±1.5) 63.1 (±0.6)
 (Yim et al. 2023) Mistral 7B 47.8 48.4 45.1 (±0.6) 50.2 (±3.0) 50.3 (±0.5)
Llama3 8B 50.5 54.2 (±0.8)# 56.2 (±0.4)#
Table 12: Comparing CriSPO with manual prompts and competitive automatic prompt engineering baseline OPRO on representative summarization benchmarks (ROUGE-1 F averaged across 3 runs). 3-shot*: 3-shot ICL with example selection. Llama3 results marked with (#) is using Claude-3 as the optimizer, due to limited context window of Llama3.
Manual Automatic Prompt Engineering
Dataset LLM 0-shot 3-shot* OPRO CriSPO CriSPO 3-shot*
CNN Claude Instant 12.5 14.8 14.3 (±0.3) 15.7 (±0.9) 17.0 (±0.2)
Claude3 Sonnet 14.4 15.4 15.1 (±0.2) 17.3 (±1.5) 16.3 (±0.5)
Mistral 7B 11.0 10.6 14.4 (±0.8) 14.3 (±0.6) 14.3 (±0.1)
Llama3 8B 14.4 15.2 (±0.4)# 16.3 (±0.9)#
MeetingBank Claude Instant 11.6 17.3 20.3 (±6.9) 23.7 (±4.7) 35.4 (±0.5)
Claude3 Sonnet 14.2 22.0 21.8 (±2.8) 32.5 (±2.2) 46.5 (±1.8)
Mistral 7B 11.5 14.8 15.4 (±2.5) 19.5 (±6.7) 16.7 (±0.9)
Llama3 8B 14.6 22.3 (±2.7)# 27.6 (±0.4)#
SAMSum Claude Instant 11.7 14.3 13.4 (±0.9) 16.9 (±2.2) 18.7 (±0.8)
Claude3 Sonnet 12.7 16.6 14.7 (±0.1) 17.1 (±1.0) 20.8 (±0.3)
Mistral 7B 10.2 14.1 13.6 (±1.4) 12.4 (±1.5) 14.2 (±1.0)
Llama3 8B 12.3 14.7 (±0.4)# 18.8 (±3.8)#
ACI-Bench Claude Instant 16.9 23.6 16.3 (±0.4) 19.7 (±0.6) 26.7 (±2.3)
Claude3 Sonnet 20.3 30.1 20.1 (±1.4) 21.4 (±0.8) 32.5 (±0.9)
Mistral 7B 17.7 19.2 17.0 (±0.1) 18.2 (±1.7) 18.7 (±0.7)
Llama3 8B 19.8 22.0 (±0.2)# 22.8 (±0.2)#
Table 13: Comparing CriSPO with manual prompts and competitive automatic prompt engineering baseline OPRO on representative summarization benchmarks (ROUGE-2 F averaged across 3 runs). 3-shot*: 3-shot ICL with example selection. Llama3 results marked with (#) is using Claude3 Sonnet as the optimizer, due to limited context window of Llama3.
Manual Automatic Prompt Engineering
Dataset LLM 0-shot 3-shot* OPRO CriSPO CriSPO 3-shot*
CNN Claude Instant 22.6 24.8 24.5 (±0.5) 26.1 (±0.4) 27.4 (±0.5)
Claude3 Sonnet 24.0 25.2 25.1 (±0.5) 27.9 (±0.9) 27.1 (±0.6)
Mistral 7B 20.4 20.1 23.0 (±1.5) 23.9 (±1.3) 24.1 (±0.7)
Llama3 8B 23.8 24.6 (±0.4)# 26.5 (±0.5)#
MeetingBank Claude Instant 20.5 25.5 29.7 (±7.4) 33.1 (±4.5) 44.4 (±0.2)
Claude3 Sonnet 22.3 29.5 32.0 (±2.8) 40.9 (±2.0) 54.1 (±1.6)
Mistral 7B 18.5 22.7 24.2 (±3.4) 29.3 (±6.5) 26.1 (±1.0)
Llama3 8B 22.6 31.5 (±3.3)# 36.8 (±0.7)#
SAMSum Claude Instant 25.6 28.8 28.7 (±1.2) 34.3 (±2.0) 36.2 (±0.2)
Claude3 Sonnet 27.0 31.3 30.1 (±1.1) 34.3 (±2.3) 38.2 (±0.5)
Mistral 7B 24.1 30.3 29.0 (±0.7) 28.4 (±2.9) 30.8 (±1.3)
Llama3 8B 27.1 30.0 (±0.5)# 35.4 (±3.4)#
ACI-Bench Claude Instant 26.1 33.5 25.5 (±1.0) 26.8 (±1.4) 35.3 (±2.3)
Claude3 Sonnet 29.3 38.6 29.5 (±1.1) 30.3 (±0.4) 41.0 (±0.6)
Mistral 7B 25.4 28.1 25.2 (±0.1) 25.6 (±1.9) 26.2 (±0.4)
Llama3 8B 27.7 29.3 (±0.6)# 29.9 (±0.5)#
Table 14: Comparing CriSPO with manual prompts and competitive automatic prompt engineering baseline OPRO on representative summarization benchmarks (ROUGE-L F averaged across 3 runs). 3-shot*: 3-shot ICL with example selection. Llama3 results marked with (#) is using Claude3 Sonnet as the optimizer, due to limited context window of Llama3.
Manual Automatic Prompt Engineering
Dataset LLM 0-shot 3-shot* OPRO CriSPO CriSPO 3-shot*
CNN Claude Instant 87.0 87.6 87.5 (±0.1) 87.2 (±0.4) 87.7 (±0.3)
Claude3 Sonnet 87.4 87.7 87.5 (±0.0) 87.8 (±0.0) 87.8 (±0.3)
Mistral 7B 85.6 85.8 87.0 (±0.1) 87.3 (±0.2) 87.3 (±0.1)
Llama3 8B 87.2 87.4 (±0.1)# 87.6 (±0.1)#
MeetingBank Claude Instant 85.0 86.0 86.7 (±1.2) 86.8 (±0.3) 89.2 (±0.1)
Claude3 Sonnet 85.4 86.9 87.1 (±0.4) 88.1 (±0.3) 90.8 (±0.3)
Mistral 7B 84.3 85.3 85.8 (±0.7) 86.2 (±0.3) 85.9 (±0.2)
Llama3 8B 85.4 86.7 (±0.6)# 87.7 (±0.2)#
SAMSum Claude Instant 89.2 89.8 89.8 (±0.2) 90.4 (±0.4) 90.7 (±0.5)
Claude3 Sonnet 89.5 90.3 89.8 (±0.4) 90.6 (±0.7) 91.3 (±0.1)
Mistral 7B 88.3 90.0 89.8 (±0.2) 89.5 (±0.6) 90.1 (±0.2)
Llama3 8B 88.7 89.9 (±0.1)# 90.7 (±0.5)#
ACI-Bench Claude Instant 85.5 88.1 85.1 (±0.3) 85.8 (±0.7) 88.1 (±0.5)
Claude3 Sonnet 85.7 89.1 85.7 (±0.5) 86.1 (±0.3) 90.0 (±0.3)
Mistral 7B 85.3 86.4 84.9 (±0.1) 85.5 (±0.8) 85.8 (±0.2)
Llama3 8B 85.1 86.1 (±0.2)# 86.6 (±0.4)#
Table 15: Comparing CriSPO with manual prompts and competitive automatic prompt engineering baseline OPRO on representative summarization benchmarks (BERTScore F averaged across 3 runs). 3-shot*: 3-shot ICL with example selection. Llama3 results marked with (#) is using Claude3 Sonnet as the optimizer, due to limited context window of Llama3.
Manual Automatic Prompt Engineering
Dataset LLM 0-shot 3-shot* OPRO CriSPO CriSPO 3-shot*
CNN Claude Instant 76.1 83.1 85.5 (±1.3) 73.9 (±12.6) 77.8 (±7.1)
(Reference: 78.7) Claude3 Sonnet 84.5 86.0 84.6 (±1.3) 84.5 (±4.9) 83.9 (±5.5)
Mistral 7B 84.9 85.2 84.5 (±5.9) 84.4 (±1.3) 86.4 (±0.5)
Llama3 8B 83.7 85.4 (±0.9)# 86.1 (±1.2)#
MeetingBank Claude Instant 72.5 70.8 59.6 (±12.8) 61.9 (±6.2) 64.0 (±2.1)
(Reference: 51.4) Claude3 Sonnet 71.9 70.8 57.5 (±3.3) 49.9 (±16.8) 70.5 (±2.1)
Mistral 7B 76.5 72.1 76.5 (±6.5) 76.5 (±2.1) 76.6 (±0.6)
Llama3 8B 72.2 71.9 (±14.9)# 63.7 (±1.3)#
SAMSum Claude Instant 85.7 86.6 84.5 (±0.9) 85.3 (±3.6) 83.9 (±1.2)
(Reference: 79.9) Claude3 Sonnet 87.9 87.2 89.5 (±0.5) 87.0 (±1.3) 84.4 (±1.5)
Mistral 7B 87.6 86.8 88.4 (±0.8) 87.7 (±0.7) 87.4 (±1.3)
Llama3 8B 88.8 88.9 (±0.7)# 87.7 (±2.5)#
ACI-Bench Claude Instant 66.7 66.3 62.3 (±2.0) 63.3 (±3.3) 65.6 (±1.1)
(Reference: 61.4) Claude3 Sonnet 70.2 67.4 69.8 (±7.8) 65.0 (±3.3) 63.8 (±0.6)
Mistral 7B 68.0 69.0 65.6 (±0.4) 67.8 (±2.3) 67.2 (±1.6)
Llama3 8B 72.5 59.4 (±1.8)# 62.3 (±1.2)#
Table 16: Comparing CriSPO with manual prompts and competitive automatic prompt engineering baseline OPRO on representative summarization benchmarks (AlignScore averaged across 3 runs). 3-shot*: 3-shot ICL with example selection. Llama3 results marked with (#) is using Claude3 Sonnet as the optimizer, due to limited context window of Llama3. Although AlignScore is among the SoTA metrics for the factual consistency, it is far from perfect. It tends to favor verbatim copies and simple paraphrases of the source document over highly abstractive summaries. As an evidence, the ground truth reference summary from the dataset gets lower AlignScore than LLM generated summary.
Appendix N Standard Deviation for QA
Manual Automatic Prompt Engineering
Dataset Claude 0-shot 64-shot OPRO CriSPO CriSPO 64-shot
Natural Question (Exact Match) Instant 34.0 33.4 8.0 (±6.6) 36.5 (±2.2) 37.8 (±1.1)
SOTA: 60.4 (Izacard et al. 2023) Sonnet 26.6 32.0 6.7 (±5.9) 38.3 (±1.6) 38.7 (±3.9)
TriviaQA (Exact Match) Instant 58.6 59.2 53.7 (±3.3) 66.3 (±1.1) 67.5 (±1.0)
SOTA: 86.1 (Touvron et al. 2023) Sonnet 58.4 65.0 41.8 (±23.9) 70.6 (±0.2) 72.1 (±0.3)
0-shot 5-shot OPRO CriSPO CriSPO 5-shot
Squad (F1) Instant 79.5 82.5 78.5 (±4.1) 87.8 (±0.5) 89.4 (±0.2)
SOTA: 95.8 (Li et al. 2020) Sonnet 76.1 83.2 76.4 (±7.4) 85.3 (±3.8) 87.9 (±2.5)
NarrativeQA (Rouge-L) Instant 64.2 67.0 59.4 (±13.2) 75.1 (±0.4) 76.1 (±0.5)
SOTA: 59.87 (Nishida et al. 2019) Sonnet 64.0 66.7 58.6 (±09.9) 76.2 (±1.6) 75.2 (±1.0)
MedMCQA (Accuracy) Instant 49.2 53.8 50.5 (±0.9) 52.3 (±2.9) 54.4 (±2.1)
SOTA: 73.7 (Nori et al. 2023) Sonnet 49.8 54.4 57.7 (±2.1) 57.9 (±0.9) 57.4 (±0.3)
Table 17: Comparing CriSPO with manual prompts and competitive automatic prompt engineering baseline OPRO on representative question answering benchmarks. *-shot: few-shot ICL with example selection.
Table 17 shows the full results for QA (Question Answering) datasets with standard deviation reported over three runs.